OpenStudy (anonymous):

Solve 3x - 5 = -2x + 5

OpenStudy (anonymous):

Solve for x: 3 x-5 = 5-2 x Move terms with x to the left hand side. Add 2 x to both sides: 3 x+2 x-5 = (2 x-2 x)+5 Look for two terms that sum to zero. 2 x-2 x = 0: 3 x+2 x-5 = 5 Add like terms in 3 x+2 x-5. 3 x+2 x = 5 x: 5 x-5 = 5 Isolate terms with x to the left hand side. Add 5 to both sides: 5 x+(5-5) = 5+5 Look for two terms that sum to zero. 5-5 = 0: 5 x = 5+5 Evaluate 5+5. 5+5 = 10: 5 x = 10 Divide both sides by a constant to simplify the equation. Divide both sides of 5 x = 10 by 5: (5 x)/5 = 10/5 Any nonzero number divided by itself is one. 5/5 = 1: x = 10/5 Reduce 10/5 to lowest terms. Start by finding the GCD of 10 and 5. The gcd of 10 and 5 is 5, so 10/5 = (5×2)/(5×1) = 5/5×2 = 2: Answer: | | x = 2

OpenStudy (anonymous):

hey he explained very well still having doubts?

OpenStudy (anonymous):

Thanks ^^